OpenStudy (anonymous):

Find the value of the variable if P is between J & K. JP = 8z - 17 ; PK = 5z + 37 ; JK = 17z - 4

OpenStudy (anonymous):

JP + PK = JK (8z - 17) + (5z + 37) = 17z - 4 13z + 20 = 17x - 4 4z = 24 z = 6

OpenStudy (anonymous):

Yes! I did it right :) Thanks!